[Ncep.nhc.nco_contacts] TSB Morning Rounds - Tue Mar 21, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Tue Mar 21 14:33:04 UTC 2017


Here is today's summary of NHC computer operations:
--- Craig

----------

TAFB:
1. An AWIPS II patch was applied this morning at 1230Z (8:30 am EDT), with  
many fixes for storm surge applications. Unfortunately, extra text windows  
now pop up when users log in. This commonly occurs with AWIPS II patches.  
Chris Mello and Matt Sardi will work to eliminate those nuisance windows.

2. Windows Sysadmin Kim Gordon has repaired and returned the situational  
awareness system to operations. The fan was clogged with "dust bunnies"  
that needed to be cleaned out.
HSU:
1. Due to several problems with 16-bit color graphics on N-AWIPS  
workstations, TSB will temporarily halt the upgrade to RHEL 6 Linux, until  
NCO devises an 8-bit color configuration that is reliable and can meet  
forecaster workflow demands.

TSB will provide NCO with a list of items (large SPFs must load quickly,  
NMAP image day/night 'Fade' bar in the lower right of the GUI must work,  
FVWM menu scripts must launch correctly, kernel panics should be very  
infrequent, etc.) that need to be addressed before the upgrade can resume.

2. AWIPS II Sysadmin Chris Mello has begun upgrading the AWIPS monitors in  
HSU and in the Fish Bowl.
Other items:

NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170321/8bf46f03/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list